OpenStudy (anonymous):

How many grams of O2(g) are needed to completely burn 27.5 g of C3H8(g)?

OpenStudy (abb0t):

burning is referring to combustion reaction, meaning, prodcts will be: \(\sf \color{red}{CO_2+ H_2O}\) So, write the equation down first and balance it, if needed. Is your first step! Can you do that.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

C3H8+5O2=3CO2+4H2O

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

thats just the balanced equation

OpenStudy (anonymous):

i need the ph solution

OpenStudy (anonymous):

you're on the wrong thread

OpenStudy (anonymous):

what

OpenStudy (anonymous):

this is my question. not yours. your post is the thread below mine on the left side of the screen

OpenStudy (anonymous):

k

OpenStudy (abb0t):

Anyways, lol. Yes, once it is balanced, now you need to use dimensional analysis to go from grams of propane \(\rightarrow\) grams of oxygen.

OpenStudy (abb0t):

So, you have 27.5 g \(\sf \color{}{C_3H_8}\) \(\sf \color{}{\times \frac{mole~C_3H_8}{grams~of~C_3H_8}}\). nOTICE how the units of grams cancel out for butane!

OpenStudy (abb0t):

To find the molas mass, look at your periodic table, add up all the values of 3carbons + 8 hydrogens. Remember that molar mass is in units of \(\sf \color{Red}{\frac{grams}{mole}}\)

OpenStudy (anonymous):

Ok, may take a bit.

OpenStudy (anonymous):

44.0962 is the molar mass

OpenStudy (anonymous):

haha no I'm not sure

OpenStudy (abb0t):

nO WAIT nevermind you're right. i was thinking butane. sorry.

OpenStudy (anonymous):

awesome

OpenStudy (abb0t):

Ok, now follow the step i provided above \(\sf \color{red}{ 27.5~ g~propane \times \frac{mole~of~propane}{44.10~g~propane} = mole~of~propane}\)

OpenStudy (abb0t):

Once you have moles of propane, you can simply multiply by 5! Since you have 5 moles of O\(_2\) and that is your final answer.

OpenStudy (anonymous):

3.18?

OpenStudy (abb0t):

Is that what you got after multiplying moles \(\times\) 5?